OpenStudy (anonymous):

please help: In rectangle PQRS, PQ = 18, PS = 14, and PR = 22.8. Diagonals PR¯¯¯¯¯ and QS¯¯¯¯¯ intersect at point T. What is the length of TQ¯¯¯¯¯? 7 9 11.4 22.8

OpenStudy (anonymous):

TQ=1/2 of QS QS=\[sqrtQR{^{2}+SR^{2}}\] =sqrt (196+324)=22.8 @bruno101 so it's D

OpenStudy (anonymous):

owww sorry TQ=1/2 of QS= 11.4 it's C @bruno101

OpenStudy (anonymous):

medal pls :) @bruno101